OpenStudy (arindameducationusc):

Without calculator,steps for finding cos inverse(8/208)?

OpenStudy (amilapsn):

\[\arccos{\frac{8}{208}}=\arccos{\frac{1}{26}}=\frac{\pi}{2}-\arcsin{\frac{1}{26}}\approx \frac{\pi}{2}-\frac{1}{26}\]because for small \(\theta\), \(\sin\theta=\theta\Rightarrow\arcsin\theta=\theta \)

OpenStudy (arindameducationusc):

Answer had to come in degrees which is 88 @amilapsn .. Please proceed again...

OpenStudy (arindameducationusc):

please do all in steps.... Till I get 88 degrees

ganeshie8 (ganeshie8):

Nice! @arindameducationusc are you saying you don't know how to convert \(\frac{\pi}{2}-\frac{1}{26}\) from radians to degrees ?

OpenStudy (amilapsn):

\[\arccos{\frac{8}{208}}\approx\frac{\pi}{2}-\frac{1}{26}=\left(\left(\frac{\pi}{2}-\frac{1}{26}\right)\times \frac{180}{\pi}\right)^o\approx88^o\]

OpenStudy (amilapsn):

Are you ok @arindameducationusc ?

OpenStudy (arindameducationusc):

It should be (pi/2-1/26)*pi/180.... @amilapsn @ganeshie8

OpenStudy (arindameducationusc):

No, I got it....my bad..... silly mistakes do happen! LOL...!!!! I got it@amilapsn

ganeshie8 (ganeshie8):

whichever you're converting to, that must be on top : radians to degrees : 180/pi degrees to radians : pi/180

OpenStudy (arindameducationusc):

Thank you. one question @ganeshie8 why did we use pi/2-arcsin(theota)?????

OpenStudy (arindameducationusc):

property? Send me a link about this property?

ganeshie8 (ganeshie8):

familiar with this identity \(\arcsin(x)+\arccos(x)=\frac{\pi}{2}\) ?

OpenStudy (arindameducationusc):

o yes, got it...! thank you.....

ganeshie8 (ganeshie8):

np :) you also need to know the small angle approximation of \(\sin\theta\) when \(\theta\) is small, we can replace \(\sin\theta \) by \(\theta\)